# SIV-Bench: A Video Benchmark for Social Interaction Understanding and Reasoning

### v1

@@ 5 Conclusion
In this work, we introduce SIV-Bench, a novel video benchmark designed to rigorously evaluate Multimodal Large Language Models’ capabilities in understanding and reasoning about complex social interactions. Featuring 2,792 diverse videos categorized by foundational social relationship models and 8,792 meticulously generated question-answer pairs, SIV-Bench structures evaluation around three core dimensions: Social Scene Understanding (SSU), Social State Reasoning (SSR), and Social Dynamics Prediction (SDP), further detailed into ten fine-grained sub-tasks. Our comprehensive experiments reveal that while leading MLLMs adeptly handle SSU, they still struggle with SDP and SSR, where Relation Inference emerges as an acute bottleneck. Moreover, our study underscores the critical impact of transcribed dialogue on comprehending complex social interactions.
Although SIV-Bench provides a strong evaluation framework, its current scale suggests room for future expansion into a larger training corpus with broader social contexts and more diverse content. The multiple-choice format also offers a basis for future extensions into interactive or generative evaluation. By identifying current model strengths and weaknesses, SIV-Bench offers timely insights and a foundation for advancing socially intelligent AI.
@@ A.1 Video Collection
(a) Grandparent-Child
(b) Parent-Child
@@
(f) Boss-Employee
(g) Coach-Player
(h) Teacher-Student
(i) Caregiver-Recipient
@@
Figure 7: Word clouds of GPT-generated keywords used for sourcing videos across 14 distinct social relationship types (e.g., (a) Grandparent-Child, (b) Parent-Child, (e) Friends, (f) Boss-Employee). Keywords were designed to capture diverse interaction scenarios, forms, and intimacy levels from platforms like YouTube and TikTok.
@@ B.1 Automatically QA Generation
B.2 Automatically QA Filtering
To ensure that the QA pairs in SIV-Bench genuinely require understanding of the video content, particularly for Subset 1, we employ a video-agnostic filtering step. Figure 12 presents the prompt used to instruct GPT for this purpose. The central directive of this prompt is for the model to select the most likely answer from the provided options using only the textual content of the question and the options themselves, without any access to the video. The prompt also enforces a strict output format, requiring only the letter of the chosen answer, and instructs the model to make a random choice if it deems the information insufficient (though the primary test is correct identification based on text alone). If a QA pair is answered correctly under this video-absent condition, it indicates that the question may be solvable through world knowledge, commonsense reasoning, or cues within the text of the QA itself, rather than necessitating specific information from the video. Such QAs are subsequently removed from the candidate pool to enhance the benchmark’s focus on video-grounded social understanding.
Figure 12: Prompt for the video-agnostic filtering of Question-Answer pairs.
To further refine the SIV-Bench dataset, particularly in selecting challenging items for Subset 1, we employ GPT-4o-mini to assign a difficulty score to each Question-Answer (QA) pair. Figure 13 presents the prompt designed for this task. It directs the model to evaluate QA difficulty on a 1 (Very Easy) to 5 (Very Hard) scale, assuming the QA pertains to a social interaction video. The prompt outlines key assessment criteria, including the subtlety of relevant cues, the complexity of reasoning involved, and the depth of social understanding needed to arrive at the correct answer. This automated difficulty scoring enables us to systematically identify and prioritize QAs that demand more sophisticated comprehension, thereby enhancing the overall rigor of our benchmark.
Figure 13: Prompt used to score the difficulty of SIV-Bench QA pairs on a 1-to-5 scale, based on defined criteria such as cue subtlety, reasoning complexity, and depth of social understanding required.
@@ B.4 QA statistics
Figure 17: Statistical analysis of SIV-Bench Question-Answer (QA) pairs. (a) average word count consistency across answer options, (b) distribution of correct answers among options, (c) overall question length distribution by word count, and (d) frequency of common first words in questions.
Appendix C Experiments Details
@@ C.1 Settings Details
For all evaluations, the specific inference parameters used for each model—such as temperature, top-p, or maximum new tokens—are adopted from their default configurations as provided within the VLMEvalKit [10] framework. This adherence to default settings aims to reflect the out-of-the-box capabilities of these models and ensure fair comparability. The experiments are conducted on two primary compute clusters. Cluster 1, utilized for evaluating the largest open-source models (Qwen2.5-VL-72B-Instruct and InternVL-78B), is equipped with an AMD EPYC 7642 48-Core Processor and 4x NVIDIA A100 GPUs. The total runtime for the reported experiments on this cluster is approximately 3 days. Cluster 2, used for the remaining models, consists of an Intel(R) Xeon(R) Platinum 8369B CPU @ 2.90GHz and 8x NVIDIA RTX 3090 GPUs. The cumulative runtime for experiments on this cluster is approximately 2 days. It should be noted that the overall research project, including preliminary testing on earlier dataset versions and exploratory experiments not included in the final results, involved a greater amount of compute time than the specific durations reported for the final benchmark evaluations.
@@
C.2 Results Details
Table 4: Comparative accuracy (%) of MLLMs on the 10 fine-grained sub-tasks in SIV-Bench (Origin), grouped by SSU, SSR, and SDP dimensions.
Models SSU SSR SDP Overall AR FER HAI EP II EI AI RI FP CP mPLUG-Owl3 62.36 68.35 59.50 70.61 62.81 65.08 61.17 60.13 56.98 74.98 63.79 LLaVA-OneVision 59.55 65.60 71.90 59.34 70.36 55.79 56.74 61.18 70.06 58.68 63.73 LLaVA-Video 60.11 63.76 71.07 78.41 64.17 61.22 56.54 60.05 65.19 61.48 63.18 Qwen2.5-VL-7B-Instruct 58.71 64.22 70.25 81.26 68.15 68.21 59.96 56.58 64.99 63.21 65.01 InternVL3-8B 64.05 72.48 67.77 82.47 70.08 70.32 61.77 53.03 60.90 70.62 66.14 Qwen2.5-VL-72B-Instruct 86.52 85.78 80.99 84.01 71.72 75.81 67.81 66.99 74.76 77.04 74.25 InternVL3-78B 82.30 81.65 82.65 82.16 73.78 73.15 63.18 61.02 72.15 74.32 72.16 o4-mini 86.80 90.83 77.69 87.26 72.42 67.46 65.80 64.73 72.05 73.83 72.30 GPT-4o 88.48 84.40 80.99 88.67 70.45 78.02 66.20 64.65 74.41 75.31 73.16 Gemini-2.0-Flash 90.73 83.03 85.12 85.21 74.11 67.03 67.20 64.41 69.87 77.04 72.15 Gemini-2.5-Flash 91.29 85.32 87.60 88.20 74.48 67.04 67.61 64.97 71.86 78.03 73.67 Gemini-2.5-Pro 93.54 91.74 89.26 88.08 75.14 76.94 69.62 67.80 72.01 79.18 76.03
Table 5: Comparative accuracy (%) of MLLMs on the 10 fine-grained sub-tasks in SIV-Bench (+sub), grouped by SSU, SSR, and SDP dimensions.
Models SSU SSR SDP Overall AR FER HAI EP II EI AI RI FP CP mPLUG-Owl3 62.60 68.55 59.69 70.83 63.03 65.30 61.36 60.33 57.17 75.22 64.01 LLaVA-OneVision 60.22 66.25 72.59 60.01 71.03 56.47 57.41 61.88 70.71 59.37 64.40 LLaVA-Video 60.65 64.30 71.61 78.98 64.75 61.75 57.12 60.63 65.76 62.07 63.75 Qwen2.5-VL-7B-Instruct 58.81 64.34 70.35 81.34 68.29 68.34 60.08 56.68 65.10 63.35 65.13 InternVL3-8B 64.21 72.60 67.90 82.65 70.27 70.44 61.89 53.17 61.01 70.75 66.28 Qwen2.5-VL-72B-Instruct 87.02 86.31 81.52 84.53 72.23 76.34 68.32 67.51 75.28 77.55 74.77 InternVL3-78B 82.87 82.21 83.20 82.69 74.31 73.66 63.73 61.57 72.69 74.90 72.70 o4-mini 86.91 90.94 77.81 87.39 72.54 67.62 65.90 64.88 72.18 74.00 72.43 GPT-4o 89.60 85.51 82.12 89.83 71.53 79.17 67.38 65.78 75.53 76.45 74.29 Gemini-2.0-Flash 91.83 84.12 86.22 86.34 75.22 68.15 68.31 65.50 70.98 78.14 73.27 Gemini-2.5-Flash 91.42 85.44 87.72 88.34 74.63 67.19 67.73 65.09 72.00 78.19 73.82 Gemini-2.5-Pro 94.02 92.25 89.77 88.52 75.61 77.43 70.09 68.25 72.43 79.66 76.50
Table 6: Comparative accuracy (%) of MLLMs on the 10 fine-grained sub-tasks in SIV-Bench (-sub), grouped by SSU, SSR, and SDP dimensions.
Models SSU SSR SDP Overall AR FER HAI EP II EI AI RI FP CP mPLUG-Owl3 61.77 67.76 58.92 70.03 62.26 64.52 60.60 59.58 56.41 74.42 63.22 LLaVA-OneVision 57.98 64.00 70.30 57.73 68.77 54.20 55.15 59.58 68.45 57.10 62.14 LLaVA-Video 58.59 62.24 69.51 76.89 62.64 59.68 55.00 58.52 63.71 59.93 61.66 Qwen2.5-VL-7B-Instruct 57.23 62.72 68.78 79.79 66.62 66.71 58.48 55.10 63.53 61.76 63.53 InternVL3-8B 63.26 71.69 66.99 81.66 69.30 69.57 60.96 52.22 60.12 69.86 65.35 Qwen2.5-VL-72B-Instruct 85.83 85.08 80.28 83.31 71.02 75.11 67.09 66.32 74.03 76.29 73.54 InternVL3-78B 81.70 81.06 82.09 81.56 73.22 72.56 62.57 60.39 71.55 73.71 71.56 o4-mini 86.10 90.11 76.97 86.57 71.73 66.75 65.09 63.99 71.35 73.12 71.59 GPT-4o 88.43 84.37 80.94 88.64 70.39 77.96 66.20 64.60 74.38 75.22 73.12 Gemini-2.0-Flash 90.24 82.50 84.63 84.73 73.55 66.55 66.73 63.89 69.35 76.52 71.65 Gemini-2.5-Flash 90.17 84.20 86.45 87.07 73.34 65.87 66.46 63.85 70.71 76.90 72.53 Gemini-2.5-Pro 92.69 90.84 88.36 87.15 74.24 76.08 68.72 66.90 71.10 78.28 75.14
C.3 Failure Cases
@@
Appendix D Broader Impact

### v2

@@ 5 Conclusion
We introduce SIV-Bench, a video benchmark for evaluating MLLMs on real-world social interaction understanding and reasoning grounded in foundational social relationship models. SIV-Bench contains 2,792 videos and 5,455 curated QA pairs, organized into three dimensions (SSU, SSR, and SDP) as well as ten fine-grained sub-tasks for diagnosis. Extensive experiments show that current MLLMs perform relatively well on SSU but struggle with high-level reasoning, with Relation Inference exhibiting systematic confusions across relation types. Subtitle ablations suggest that textual signals can affect performance in a task-dependent manner, particularly for complex inference cases. On a challenging hard subset, we further observe a substantial gap to human performance, reinforcing the difficulty of human-aligned social reasoning. We hope SIV-Bench will serve as a unified testbed to track progress and drive more robust social intelligence in MLLMs.
Limitations
While SIV-Bench offers a structured diagnostic benchmark for social interaction understanding, the current size and coverage still leave room to expand toward broader social settings, cultures, and interaction styles. Moreover, our primary evaluation remains multiple-choice, which simplifies scoring but does not fully capture open-ended grounding, generation, or interactive behaviors. We include SIV-Bench-Hard with brief explanations to partially address this gap, yet it is relatively small and still constrained in evaluation scope. Finally, social interpretation can be subjective and culturally contingent, and some clips may admit multiple plausible readings, suggesting the need for future work on more calibrated human protocols and richer evaluation formats.
Ethical Considerations
SIV-Bench is built from publicly available videos (e.g., TikTok and YouTube), and we carefully consider licensing and privacy risks when creating and releasing the benchmark. To reduce copyright and platform-policy concerns, we will not redistribute raw video files; instead, we will release annotations along with video identifiers (e.g., URLs and timestamps) and provide a script for users to retrieve content directly from the original platforms, following common practice in prior video benchmarks.
Because the videos may contain identifiable individuals, we applied manual filtering to exclude content that appears overly private or sensitive, and we encourage responsible downstream use. Our human annotation procedures were designed to be minimal-risk, with annotators compensated and handled anonymously. Finally, while SIV-Bench is intended to support transparent evaluation of MLLMs’ social understanding, we acknowledge the dual-use nature of social inference technologies and recommend that users avoid deployment in high-stakes settings and follow applicable privacy and platform guidelines when using the benchmark.
@@ A.1 Video Collection
(a) Grandparent–Child
(b) Parent–Child
@@
(f) Boss–Employee
(g) Coach–Player
(h) Teacher–Student
(i) Caregiver–Recipient
@@
Figure 8: Word clouds of GPT-generated keywords used for sourcing videos across 14 distinct social relationship types. Keywords capture diverse interaction scenarios, roles, and intimacy levels from platforms such as YouTube and TikTok.
@@ B.2 Human Annotation
Quality Control and Inter-Annotator Agreement. To quantitatively address label quality, we compute the Inter-Annotator Agreement (IAA) using Fleiss’ Kappa. For the initial set of approximately 27,000 QA pairs generated by our three LLM experts, the resulting Fleiss’ Kappa is 0.52, indicating "moderate" agreement. For the questions that the LLMs disagreed, we measure the agreement between our human annotators, yielding a Fleiss’ Kappa of 0.68. According to established standards (Landis and Koch, 1977), this value represents "substantial" agreement. Achieving this on the most difficult portion of our dataset confirms that humans can establish a reliable ground truth and justifies our methodology of retaining only those questions on which all assigned annotators unanimously agreed for the final benchmark.
@@ B.3 QA statistics
Figure 17: Statistical analysis of SIV-Bench question–answer (QA) pairs. (a) Average word count consistency across answer options, (b) distribution of correct answers among options, (c) question length distribution by word count, and (d) frequency of common first words in questions.
B.4 Quality and Diversity Analysis
To further validate the quality of SIV-Bench, we conduct two analyses to ensure our benchmark encourages genuine comprehension over fitting to superficial patterns.
B.4.1 Analysis of Template Pattern Exploitation
SIV-Bench organizes tasks into SSU, SSR, and SDP to evaluate distinct dimensions of social cognition. To directly test whether models exploit surface-level cues, we analyze the properties of questions that a representative model (Gemini-2.0-Flash) answers correctly versus incorrectly. As shown in Table 5, while statistically significant differences in length and word count exist, the small effect sizes (Cohen’s d) indicate no meaningful structural separation between the two groups. Furthermore, Table 6 shows a high cosine similarity between the embeddings of the correct and incorrect QA sets, along with similar internal distributions (intra-similarity and variance). This close alignment suggests that model performance is unlikely to rely on superficial statistical or semantic patterns.
Table 5: Length and word count comparison between correctly and incorrectly answered questions.
Metric Mean Length (chars) Word Count
Correct 251.28 ± 93.88 42.13 ± 15.50
Wrong 228.96 ± 94.49 38.21 ± 15.78
t-stat 9.05 9.54
p-value 2.36e-19 2.63e-21
Cohen’s d 0.2373 0.2516
Table 6: Semantic similarity analysis between correctly and incorrectly answered question sets.
Metric Value
Cosine Similarity (TF-IDF) 0.9765
Cosine Similarity (SentenceTransformer) 0.9759
Correct Intra-similarity 0.1689
Wrong Intra-similarity 0.1711
Correct Embedding Variance 0.002164
Wrong Embedding Variance 0.002157
B.4.2 Linguistic Diversity Analysis
The majority of questions in SIV-Bench are generated by large language models or written by human annotators, rather than using rigid templates. To objectively measure our benchmark’s linguistic diversity, we compared it against several prominent video QA benchmarks using two standard metrics: Mean Semantic Distance (average pairwise cosine distance of sentence embeddings) and Vector Variance (average variance across embedding dimensions). As shown in Table 7, SIV-Bench exhibits high semantic diversity, ranking among the top benchmarks. These results support that SIV-Bench’s questions are varied and not limited to shallow templates, thereby promoting genuine semantic understanding.
Table 7: Semantic diversity comparison across video QA benchmarks.
Benchmark Mean Dist. $\uparrow$ Variance $\uparrow$
SIV-Bench 0.8321 0.0022
Video-Bench 0.8811 0.0023
Perception_Test 0.7677 0.0020
VideoVista 0.7604 0.0020
Social-IQ 2.0 0.7260 0.0019
MVBench 0.7124 0.0019
EgoSchema 0.5411 0.0014
Appendix C Experimental Details
@@ C.1 Settings Details
For all evaluations, the specific inference parameters used for each model—such as temperature, top-p, or maximum new tokens—are adopted from their default configurations as provided within the VLMEvalKit (Duan et al., 2024) framework. This adherence to default settings aims to reflect the out-of-the-box capabilities of these models and ensure fair comparability. The experiments are conducted on two primary compute clusters. Cluster 1, utilized for evaluating the largest open-source models (Qwen2.5-VL-72B-Instruct and InternVL-78B), is equipped with an AMD EPYC 7642 48-Core Processor and 4x NVIDIA A100 GPUs. The total runtime for the reported experiments on this cluster is approximately 3 days. Cluster 2, used for the remaining models, consists of an Intel(R) Xeon(R) Platinum 8369B CPU @ 2.90GHz and 8x NVIDIA RTX 3090 GPUs. The cumulative runtime for experiments on this cluster is approximately 2 days. It should be noted that the overall research project, including preliminary testing on earlier dataset versions and exploratory experiments not included in the final results, involved a greater amount of compute time than the specific durations reported for the final benchmark evaluations.
@@
C.2 Failure Cases
@@
C.3 Text-Only Control for Shortcut Analysis
To further quantify the extent to which SIV-Bench can be solved from textual priors alone, we introduce a text-only control that removes all visual and audio information and provides models with only the question and the five answer options. This setting is designed as a direct probe of textual shortcut solvability: if a substantial portion of the benchmark could be solved without access to the video, then models should retain strong performance under this control.
Setup.
We evaluate five representative MLLMs under the text-only setting, including both closed- and open-source systems. In this control, models receive the same question and candidate options as in the standard evaluation, but no video frames, no audio, and no subtitle input are provided. Since each question in SIV-Bench has five candidate answers, the random-guessing baseline is $20\%$. We report accuracy on the full benchmark and on the three task categories, namely Social Scene Understanding (SSU), Social State Reasoning (SSR), and Social Dynamics Prediction (SDP).
Results.
As shown in Table 8, text-only performance remains close to the random baseline across all five models, and the ordering of models is notably less stable than in the standard multimodal setting. This suggests that, once visual and audio evidence are removed, models largely lose the consistent advantage observed in the full benchmark. We also observe that SSU does not consistently become easier in the text-only setting, which is expected since many SSU questions require direct grounding in visual content. Meanwhile, some SSR and SDP items remain marginally solvable from textual priors alone, but the gains are still small and far below the standard multimodal setting. Moreover, the proportion of questions answered correctly by all five models, or by at least four out of five models, stays very low. These findings indicate that SIV-Bench contains relatively few questions that can be reliably solved from textual priors alone, and they further support our claim that strong performance on the more reasoning-intensive portions of SIV-Bench, especially SSR and SDP, cannot be adequately explained by text-only shortcuts.
Model / Indicator Overall SSU SSR SDP $\Delta$ vs Rand.
Gemini-2.5-Pro (text-only) 24.0 23.0 22.0 26.0 +4.0
GPT-4o (text-only) 23.0 21.0 24.0 22.0 +3.0
Qwen2.5-VL-7B (text-only) 22.0 24.0 20.0 21.0 +2.0
InternVL3-8B (text-only) 25.0 20.0 23.0 29.0 +5.0
LLaVA-Video (text-only) 21.0 19.0 22.0 23.0 +1.0
Random (5-way) 20.0 20.0 20.0 20.0 +0.0
All 5/5 models correct (%) 0.7 0.9 0.6 0.8
At least 4/5 models correct (%) 4.2 5.8 3.6 4.9
Table 8: Text-only control and dataset-level shortcut indicators on SIV-Bench. In the text-only setting, models receive only the question and the five answer options, without any visual, audio, or subtitle input. The results remain close to random guessing, and the low cross-model agreement further suggests that only a small fraction of questions can be solved from textual priors alone.
C.4 Analysis of Chain-of-Thought Prompting
To investigate the impact of explicit reasoning on model performance, we conducted a preliminary experiment using a Chain-of-Thought (CoT) prompting strategy. We prepended the instruction "Let’s think step by step. First, output your reasoning process, and then output the final answer." to our standard evaluation prompt. The overall accuracy on the origin videos, with and without CoT, is presented in Table 9.
The results indicate that applying a generic CoT prompt did not yield significant performance improvements for most models. For several smaller open-source models (e.g., mPLUG-Owl3, LLaVA-Video), it resulted in a notable performance decrease. We observed that this is often because these models struggle to consistently adhere to the more complex two-stage output format (i.e., providing reasoning before the final answer in the required format), leading to failures in our answer parsing logic.
The primary goal of SIV-Bench is to establish a fair, consistent, and reproducible evaluation of baseline model capabilities. Our standardized prompting strategy, which aligns with widely used toolkits like VLMEvalKit, ensures this fairness. While techniques like CoT are powerful for eliciting maximum performance from certain capable models (e.g., Gemini-2.5-Pro), introducing them as a default can create a confounding variable. Such a setup might shift the evaluation from testing inherent social reasoning to testing complex instruction-following abilities. Therefore, our main experiments use a direct-answering prompt to maintain a level playing field. Nonetheless, these findings suggest that developing more specialized reasoning methods tailored to social intelligence is a valuable direction for future work that builds upon this benchmark.
Table 9: Comparison of Overall Accuracy (%) on the origin videos with and without Chain-of-Thought (CoT) prompting.
Model Acc (Origin) Acc (CoT)
mPLUG-Owl3 42.06 41.79
LLaVA-OneVision 41.97 41.22
LLaVA-Video 41.09 40.13
Qwen2.5-VL-7B-Instruct 44.02 44.76
InternVL-8B 45.82 44.99
Qwen2.5-VL-72B-Instruct 58.80 59.13
InternVL-78B 55.46 55.25
o4-mini 55.68 55.79
GPT-4o 58.02 57.91
Gemini-2.0-Flash 56.40 56.53
Gemini-2.5-Flash 57.87 57.71
Gemini-2.5-Pro 61.65 61.77
C.5 Statistical Significance Analysis
To validate the reliability of our comparative claims, we conducted McNemar’s tests on the full dataset ($N=5,455$). This paired non-parametric test is appropriate for comparing the performance of two classifiers on the same dataset.
Model Ranking.
We verified the leadership of our SOTA model. The performance difference between the top-performing Gemini-2.5-Pro (61.65%) and the second-best model Qwen2.5-VL-72B (58.80%) is highly statistically significant ($p<0.001$), confirming the robustness of the leaderboard rankings.
Task Difficulty.
We confirmed that the performance stratifications across our three core dimensions are not due to chance. For Gemini-2.5-Pro, the performance gaps between SSU (85.07%) and SDP (60.45%), as well as between SDP and SSR (54.30%), are all highly statistically significant ($p<0.001$).
Subtitle Influence.
We performed significance testing on the subtitle conditions (Table 4). Our analysis reveals that the minor Overall improvement from adding subtitles (‘+sub‘) is not statistically significant ($p=0.12$). However, the impact is task-dependent: the negative impact of removing text (‘-sub‘) is statistically significant for the SSR task ($p<0.05$), and the benefit of added subtitles (‘+sub‘) is significant for the SDP task ($p<0.05$).
Appendix D SIV-Bench-Hard Details
To rigorously evaluate the upper limits of current MLLMs, establish a robust human baseline, and probe the reasoning processes beyond simple accuracy, we curated and analyzed a challenging subset of our dataset, termed SIV-Bench-Hard. This section details the setup, human performance, and a multi-dimensional analysis of model reasoning quality on this subset.
D.1 Experimental Setup
Dataset Curation. We selected a subset of 200 questions from the human-generated portion of SIV-Bench. These questions were specifically chosen for their complexity and reliance on deep social understanding, filtering out items that could be solved via superficial visual cues.
Task Definition. Unlike the standard multiple-choice evaluation, this study required both human annotators and MLLMs to provide: (1) the selected answer option, and (2) a free-text reasoning explanation justifying their choice. The prompt is shown in Figure 22 and 23. This allows for a deeper examination of the cognitive process.
Figure 22: Model prompt for SIV-Bench-Hard.
Figure 23: LLM-judge prompt for scoring explanations in SIV-Bench-Hard.
Participants. We recruited 3 independent human annotators to perform this task to establish a human baseline. We evaluated a suite of state-of-the-art MLLMs, including Gemini-3-Pro, GPT-5.1, Gemini-2.5-Pro, Gemini-2.5-Flash, Qwen2.5-VL-7B, and GPT-4o-mini.
D.2 Quantitative Analysis of Reasoning
To quantify the divergence observed in the qualitative scores, we performed an embedding-based analysis. We encoded all reasoning texts (both human and model) using the paraphrase-multilingual-MiniLM-L12-v2 model.
Distinguishability. We trained a Random Forest classifier to distinguish between human and model explanations based on their embeddings. The classifier achieved an accuracy of 92.17% (compared to a 50% random baseline). This high classification accuracy indicates that the latent semantic features of model reasoning are fundamentally distinct from those of humans.
Statistical Significance. We further validated this difference using Mann-Whitney U tests across the embedding dimensions. The tests confirmed that the distributions of human and model embeddings are significantly different ($p<0.05$) on 8 out of 10 principal dimensions. Collectively, these results provide quantitative evidence that current MLLMs, despite their linguistic fluency, employ reasoning processes that are statistically distinguishable from human social cognition.
The SIV-Bench-Hard subset, along with the human reasoning annotations and analysis code, will be released to facilitate future research into bridging this gap.
D.3 Correlation Between Answer Correctness and Reasoning Fidelity
To rigorously verify that model performance is grounded in genuine social reasoning rather than superficial visual shortcuts (e.g., background cues or object co-occurrence), we analyze the relationship between the correctness of a model’s answer and the semantic quality of its reasoning trace. We hypothesize that if models were merely "guessing" via shortcuts, their generated explanations would lack alignment with human cognitive processes even when they fortuitously select the correct option.
We perform an embedding-based analysis on the SIV-Bench-Hard subset using the paraphrase-multilingual-MiniLM-L12-v2 model. For each of the six evaluated models, we calculate the cosine similarity between the model’s generated reasoning and the human expert ground truth. The results are stratified based on whether the model answers the multiple-choice question correctly or incorrectly.
Figure 24: Box plots illustrating the distribution of semantic similarity scores between model reasoning and human ground truth, stratified by answer correctness. We observe a consistent trend where reasoning traces for correct answers exhibit significantly higher alignment with human social cognition compared to incorrect answers. Statistical significance is denoted by * ($p<0.05$), ** ($p<0.01$), and *** ($p<0.001$).
Table 10: Quantitative comparison of reasoning similarity to human ground truth between correct and incorrect responses. The Gap ($\Delta$) represents the increase in similarity when the model answers correctly. Significance is calculated using the Mann-Whitney U Test.
Model Similarity (Correct) Similarity (Incorrect) Gap ($\Delta$) Significance ($p$-value)
GPT-5.1 0.572 0.506 +0.066 << 0.001 (***)
Qwen2.5-VL-7B 0.535 0.474 +0.061 << 0.01 (**)
GPT-4o-mini 0.571 0.513 +0.058 << 0.01 (**)
Gemini-2.5-Pro 0.541 0.498 +0.043 << 0.05 (*)
Gemini-2.5-Flash 0.537 0.504 +0.033 << 0.05 (*)
Gemini-3-Pro 0.527 0.494 +0.033 0.067 (n.s.)
As illustrated in Figure 24 and summarized in Table 10, we observe a statistically significant positive gap in similarity scores across 5 out of the 6 models. Notably, GPT-5.1 demonstrates the strongest effect, with a similarity gap of 0.066 ($p<0.001$).
These findings provide empirical evidence that correctness in SIV-Bench is strongly correlated with human-like social reasoning. The significant degradation in reasoning alignment during failure cases suggests that models do not rely on "guessing" via superficial cues; rather, successful performance necessitates a cognitive process that mirrors human social understanding.
Appendix E Broader Impact